# A Video is Worth 10,000 Words: Training and Benchmarking with Diverse Captions for Better Long Video Retrieval

### v1

@@ 7.2 Automatic Analysis
Metric
Source Short Medium Full Length Primary Secondary University S and P S and S S and U
$\Delta$ Nouns
-11.77 -4.23 1.49 -1.40 3.75 11.14 -14.35 -13.18 -12.36
$\Delta$ Verbs
-2.60 0.90 4.32 1.96 7.63 11.71 -3.07 -2.34 -1.86
Word Count
155.40 36.06 76.30 105.43 129.52 136.58 154.13 28.94 31.85 34.83
Word Length
4.66 5.00 4.96 5.18 4.79 5.25 5.74 4.66 4.90 5.12
@@
Metric
Source Short Medium Full Length Primary Secondary University S and P S and S S and U
$\Delta$ Nouns
-27.25 -20.38 -14.81 -14.83 -9.26 -2.55 -32.21 -31.16 -29.27
$\Delta$ Verbs
-12.10 -7.90 -4.46 -3.04 1.26 4.56 -14.11 -13.57 -12.83
Word Count
207.86 53.41 86.69 114.55 150.97 164.26 181.92 34.81 37.28 43.10
Word Length
5.47 5.89 5.72 5.79 5.27 5.66 6.02 5.37 5.73 5.98
@@ 7.3 Annotator Analysis
For an example survey, please refer to the attached material.
@@ 8 Ablations
Table 15: Projection ratio and text to text ratio ablations.
Hyperparameter ActivityNet
$\eta$ $\alpha_{\text{proj}}$ $\alpha_{\text{t2t}}$ Full Short Long
0.75 0.1 0.1 59.9 33.5 56.2
0.75 0.1 0.0 59.3 33.2 56.5
0.75 0.1 0.2 59.6 33.3 56.5
0.75 0.1 0.5 59.7 33.4 56.7
0.75 0.0 0.1 59.9 33.4 56.5
0.75 0.2 0.1 59.1 33.5 55.7
0.75 0.5 0.1 59.7 33.6 56.5
0.0 0.1 0.1 59.4 31.9 55.8
0.25 0.1 0.1 60.1 33.2 56.6
0.5 0.1 0.1 59.4 33.3 56.5
1.0 0.1 0.1 59.3 33.5 56.6

### v2

@@ 8.2 Automatic Analysis
Metric Source Short Medium Full Length Elementary Intermediate University S and P S and S S and U
$\Delta$ Nouns -11.77 -4.23 1.49 -1.40 3.75 11.14 -14.35 -13.18 -12.36
$\Delta$ Verbs -2.60 0.90 4.32 1.96 7.63 11.71 -3.07 -2.34 -1.86
Word Count 155.40 36.06 76.30 105.43 129.52 136.58 154.13 28.94 31.85 34.83
Word Length 4.66 5.00 4.96 5.18 4.79 5.25 5.74 4.66 4.90 5.12
@@
Metric Source Short Medium Full Length Elementary Intermediate University S and P S and S S and U
$\Delta$ Nouns -27.25 -20.38 -14.81 -14.83 -9.26 -2.55 -32.21 -31.16 -29.27
$\Delta$ Verbs -12.10 -7.90 -4.46 -3.04 1.26 4.56 -14.11 -13.57 -12.83
Word Count 207.86 53.41 86.69 114.55 150.97 164.26 181.92 34.81 37.28 43.10
Word Length 5.47 5.89 5.72 5.79 5.27 5.66 6.02 5.37 5.73 5.98
@@ 8.3 Annotator Analysis
For our sample, we recruited 15 individuals, all of whom had at least a bachelor’s degree. Individuals spent between 10 and 20 minutes to answer the 15 questions on their assigned survey. For an example survey, please refer to the attached material.
@@ 9 Ablations
Table 15: Mixing ratio ablations.
ActivityNet
$\eta$ Full Short Long
0.0 59.4 31.9 55.8
0.25 60.1 33.2 56.6
0.5 59.4 33.3 56.5
0.75 59.9 33.5 56.2
1.0 59.3 33.5 56.6
@@
10 Miscellaneous
10.1 Hallucination Prevalence Results
In Table 6 we give results computed in two ways, as the percentage of all votes which belong to a given category (“Total”) and by determining the majority label for each word, then computing percentages (“Majority”). To further clarify this computation, consider the following example, with 3 voters and 3 words. For the first word, 2 voters select matches, 1 selects unsure. For the second word, all 3 voters select unsure. For the third word, 3 select different. Since there were 3 votes for different, 4 for unsure, and 2 for matches, the percentages for total would be 33.33%, 44.44%, and 22.22% respectively. For majority, since the first was majority matches, second was majority unsure, and third was majority different, these would be 33.33% each.
10.2 Training-time Improvement Details
Figure 5: We perform contrastive finetuning for retrieval with video-caption pairs.  We propose efficient sampling of our 10k text captions for data augmentation, where we compute standard contrastive loss, but each caption is sampled  randomly from the 10k captions for a given video, according to a mixing ratio, $\eta$.
First, we show an illustration of our data sampling approach, as a visual aid, in Figure 5.
Figure 6: We measure how much our data augmentation helps in the data constrained regime, training only with the indicated amounts of data, and performing retrieval with the resulting trained models. We show that finetuning COSA with 10k data (ours) is superior to generic COSA finetuning (C) for ActivityNet10k.
Since part of our contribution is a data augmentation strategy, we also evaluate its performance by finetuning with different fractions of the original ActivityNet data in Figure 6. Notice that the absolute differences in recall between training with 10k data and training without remain consistent for all amounts of training data. For training on short captions the difference is around a 3% improvement while for long captions it is around 2%. By training with synthetic data, we achieve the same performance with less manually annotated data.
Table 16: LFVILA COSA finetuning.  Results improve with 10k finetuning.
Finetune Method All Short Long Partial
R@1 Avg. R R@1 Avg. R R@1 Avg. R R@1 Avg. R
Domain 77.3 86.9 65.2 78.4 90.2 95.9 73.8 84.9
Ours 85.2 92.6 78.2 89.2 95.3 98.2 73.0 83.9
We also show that our findings hold when finetuning on other datasets, such as LF-VILA (Table 16).
10.3 Inference-time Improvement Details
For our ensembles in Section 4.2, for the sake of simplicity, for synthetic captions we choose the ‘l’ and ‘l+i’ captions, since we find that ‘l+e’ and ‘l+u’ have higher tendency to either reduce information (for ‘l+e’) or else infer unnecessary detail (for ‘l+u’). Sampling short and medium length captions is less effective in this regime due to the information loss. Introducing such ambiguity into the retrievals would be counterproductive. To actually perform the retrieval, we compute the standard text-video similarity matrix, as well as a separate text-video similarity matrix for each type of 10k caption (‘l’ and ‘l+i’). We then add these together, giving $50\%$ weight to the standard text-video matrix, and equal weight to the remaining 2 matrices.
10.4 Information Loss and Uniqueness Details
Figure 7: We measure the number of nouns and retrieval uniqueness for short caption retrieval, and find that the highest ranks correlate with captions that have lost their nouns and unique information.
Figure 8: We measure the number of nouns and retrieval uniqueness for short caption retrieval, and find that the highest ranks correlate with captions that have lost their similarity with the source caption.
We realize the length is not a perfect measure of information. In fact, part of the motivation of this work is that captions can be quite short but very information-dense. So, we compute information loss is 3 ways. First, we use short length divided by standard length, as given in the main paper in Figure 3. Second, we use spaCy to count entities in the short and standard captions, dividing the number in the short by the number from the source standard caption in Figure 8. Third, we get the word2vec embeddings for the entities in the short and standard captions, and compute the cosine similarities between all entities. We choose the best matches for the entries in the short caption, and sum the similarities, then divide by the number of entries in the short caption. Hence we use similarity between bags of words as our proxy for how much the information in the short caption overlaps the information in the standard caption, with results in Figure 8. These two alternatives confirm the findings from using length, so we opt to use length in the main paper since it is simpler.
To calculate uniqueness, we take the similarity score defined above (greedy matching of cosine similarities for word2vec embeddings of entities). We additionally compute the similarity between the short caption and the standard captions for the top 5 retrieved videos, as retrieved using the short caption, not including the standard caption for the matching video. That is, if the matching video is in the top 5 retrievals, we exclude it and additionally consider the standard caption for the video retrieved at rank 6. We average the similarities between the short caption and these 5 standard captions, and subtract it from the similarity between short and source (matching) standard caption, for a uniqueness score. This “uniqueness” score provides the color in Figure 3, Figure 8, and Figure 8.